# Dark Matter and Electroweak Baryogenesis with Spontaneous $CP$ Violation in the Early Universe

### v1

@@ 1 Introduction
In the EWBG framework, a strong first-order electroweak phase transition (SFOEWPT) provides the requisite departure from equilibrium via expanding bubbles of the broken phase. $CP$-violating interactions in (or on) the bubble wall generate chiral charge densities that diffuse into the unbroken phase, where baryon-number-violating electroweak sphaleron processes convert them into a net baryon asymmetry Klinkhamer:1984di ; Gavela:1994dt ; Huet:1994jb . Once the bubbles expand and fill the universe, the sphaleron rate in the broken phase becomes exponentially suppressed, thereby freezing in the generated asymmetry. However, the SM itself cannot realize EWBG. Lattice and perturbative analyses show that, for the observed SM-like Higgs ($h$) mass $m_{h}\simeq 125~\mathrm{GeV}$ ATLAS:2012yve ; CMS:2012qbp , the electroweak phase transition (EWPT) is a crossover rather than first-order Kajantie:1996mn ; Csikor:1998eu ; Aoki:1999fi ; the Higgs quartic coupling $\lambda_{h}$ suppresses the thermal cubic term, eliminating the barrier needed for an SFOEWPT Bochkarev:1987wf ; Kajantie:1995kf . In addition, $CP$ violation from the CKM phase is far too small to account for the observed asymmetry Shaposhnikov:1987tw ; Farrar:1993hn ; Gavela:1993ts ; Konstandin:2003dx ; Kapusta:2006pm . Consequently, achieving a strongly first-order transition and sufficient $CP$ violation for EWBG requires BSM physics. This mechanism has motivated extensive study of the EWPT and its realizations in BSM theories Cohen:1993nk ; Rubakov:1996vz ; Trodden:1998ym ; Riotto:1998bt ; Cline:2006ts ; Morrissey:2012db ; White:2016nbo ; Cline:2017jvp ; Wagner:2023vqw ; vandeVis:2025efm . Despite its appeal, no conclusive evidence for new electroweak-scale excitations has yet emerged. Furthermore, any additional $CP$-violating interactions introduced at the electroweak scale are strongly constrained by precision measurements of electric dipole moments (EDM) of the electron ACME:2018yjb , neutron Abel:2020pzs , and heavy atoms such as mercury Griffith:2009zz . A viable way to reconcile these limits with successful EWBG is through spontaneous $CP$ violation that occurs only during the EWPT. In this picture, the temporary breaking of $CP$ in the early universe supplies the baryogenesis source while the symmetry is restored at zero temperature, naturally evading the present EDM constraints.
@@ 2.1 The Inelastic Higgs-Portal Dark Matter Model
We add quartic self-interactions for the singlet fields $\phi_{1}$ and $\phi_{2}$, parameterized by $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{12}$. In the special case $\lambda_{1}=\lambda_{2}=\lambda_{12}$, the scalar sector retains an approximate $U(1)$ symmetry. More generally, allowing these couplings to differ explicitly breaks this symmetry. In the limits of $f_{1}\sim 0$ and $\sin 2\theta=0$, the coupling relations given in Eq. (13) to Eq. (15) and the mass relation are reduced to,
@@ 2.2.1 Quantum Corrections at Zero Temperature
One-loop corrections shift both the location of the electroweak vacuum and the scalar masses and mixings. To preserve the physical mass spectrum and vacuum expectation values at one loop, counterterms $V_{\rm CT}$ are introduced, parameterized by
@@ 2.2.2 Thermal Corrections
where the thermal functions $J_{B}$ and $J_{F}$ correspond to bosons and fermions respectively:
@@ 3 Dark Matter Phenomenology
To assess the relative impact of the channels, we show $\Omega_{\phi_{1}}h^{2}$ versus $m_{\phi_{1}}$ for benchmark choices of $(\Delta m,\,g,\,f_{2})$. The coupling $f_{1}$ is fixed to a representative value $10^{-5}$, chosen sufficiently small to keep elastic Higgs-mediated scattering safely below current direct-detection bounds Hooper:2025fda ; LZCollaboration:2024lux . The horizontal dashed line marks the Planck value $\Omega_{\rm DM}h^{2}\simeq 0.12$ Planck:2018vyg ; vertical dotted lines at $m_{\phi_{1}}\simeq m_{h}/2$ and $m_{\phi_{1}}\simeq m_{h}$ indicate the Higgs resonance and the $hh$ threshold. Red/blue (green/black) curves correspond to $\Delta m=10$ GeV (0.1 GeV).
@@ 4 FOEWPT-favored parameter space
Figure 4: [Left] Scan in the $m_{\phi_{2}}$$f_{2}$ plane showing points that realize an FOEWPT. We vary $m_{\phi_{2}}\in[100,250]~\mathrm{GeV}$ and $\Delta m\in[10,20]~\mathrm{GeV}$, while holding $\lambda_{1,2,12}\sim 0.5$ and $g\sim 0.05$ fixed. [Right] Peak amplitude $\left(\Omega_{\mathrm{GW}}h^{2}\right)_{\mathrm{peak}}$ versus peak frequency $f_{\mathrm{peak}}$ of the predicted stochastic GW signal from FOEWPT-allowed points. Sensitivity curves for prospective GW observatories LISA, Taiji, ALIA, BBO, and UDECIGO are overlaid for comparison. In both panels, the color scale encodes the transition strength $\xi_{n}\equiv v_{n}/T_{n}$.
@@ 5 Spontaneous $CP$ violation and the generated baryon asymmetry
Benchmark point.- To illustrate our scenario, we specify a benchmark (Tab. 1) in which the DM mass is $\sim\!133~\mathrm{GeV}$ and the mass splitting is large enough that the relic density is set primarily by $\phi_{1}\phi_{1}\!\to hh$ via $\phi_{2}$ exchange annihilation channel. This choice yields a present-day di-Higgs indirect-detection cross section of $\langle\sigma v\rangle_{\rm ID}^{hh}\simeq 1.9\times 10^{-26}~\mathrm{cm^{3}\,s^{-1}}$, consistent with interpretations of the bright, statistically significant GeV excess from the Galactic Center for these masses Goodenough:2009gk ; Hooper:2010mq ; Hooper:2011ti ; Abazajian:2012pn ; Hooper:2013rwa ; Gordon:2013vta ; Daylan:2014rsa ; Calore:2014xka ; Zhou:2014lva ; Fermi-LAT:2015sau ; Fermi-LAT:2017opo ; Cholis:2021rpp ; DiMauro:2021raz ; Hooper:2019xss ; Hooper:2025fda . The elastic $\phi_{1}$$h$ coupling is chosen such that the spin–independent scattering rate ($\sigma_{\rm{DD}}^{\rm SI}$) lies below the “neutrino fog” Billard:2013qya .
This point also realizes the two-step thermal history central to our EWBG study. A second-order transition at $T_{n}=335~\mathrm{GeV}$ first breaks the $Z_{2}$ symmetry along the singlet directions so that $\langle\phi_{1,2}\rangle\neq 0$. At a lower temperature, $T_{n}=95~\mathrm{GeV}$, a first-order transition occurs in which electroweak symmetry is broken while the $Z_{2}$ is restored by driving the singlet $\it vevs$} back to zero. The field values at the false and true minima of the second transition are listed in Tab. 1. The corresponding bubble wall profiles $h(z)$, $\phi_{1}(z)$ and $\phi_{2}(z)$ are shown in the left plot of Fig. 5. As discussed earlier, it can be seen that all three fields remain nonzero in certain region (within the bubble wall region), where the $CP$-violating source built. In this benchmark scenario $L_{w}T_{n}=3.5$ and we set $v_{w}=0.1$. The generated BAU tracks the effective left-handed baryon chemical potential $\mu_{B_{L}}(z)$ defined in Eq. (76), whose profile for this point is shown in the right plot of Fig. 5. To match the observed value of BAU, we vary the cutoff scale and fix it at $\Lambda\simeq 800~\mathrm{GeV}$ for this benchmark point. A deep scan in the surrounding parameter space can be performed, which is expected to preserve the qualitative features of the benchmark while achieving the observed BAU with somewhat larger values of $\Lambda$.
@@ Acknowledgments
The author gratefully acknowledges Dan Hooper, Gordan Krnjaic, Tong Ou, Duncan Rocha and Carlos Wagner for many insightful discussions. This work is supported by the U.S. Department of Energy under contract No. DEAC02-06CH11357 at Argonne National Laboratory. The author further acknowledges the hospitality of the University of Chicago, Fermilab, the Aspen Center for Physics, and the Perimeter Institute for Theoretical Physics, where a significant part of this research was carried out. He is especially indebted to Carlos Wagner for hosting him at Perimeter Institute.

### v2

@@ 1 Introduction
In the EWBG framework, a strong first-order electroweak phase transition (SFOEWPT) provides the requisite departure from equilibrium via expanding bubbles of the broken phase. $CP$-violating interactions in (or on) the bubble wall generate chiral charge densities that diffuse into the unbroken phase, where baryon-number-violating electroweak sphaleron processes convert them into a net baryon asymmetry Klinkhamer:1984di ; Gavela:1994dt ; Huet:1994jb . Once the bubbles expand and fill the universe, the sphaleron rate in the broken phase becomes exponentially suppressed, thereby freezing in the generated asymmetry. However, the SM itself cannot realize EWBG. Lattice and perturbative analyses show that, for the observed SM-like Higgs ($h$) mass $m_{h}\simeq 125~\mathrm{GeV}$ ATLAS:2012yve ; CMS:2012qbp , the electroweak phase transition (EWPT) is a crossover rather than first-order Kajantie:1996mn ; Csikor:1998eu ; Aoki:1999fi ; the Higgs quartic coupling $\lambda_{h}$ suppresses the thermal cubic term, eliminating the barrier needed for an SFOEWPT Bochkarev:1987wf ; Kajantie:1995kf . In addition, $CP$ violation from the Cabibbo–Kobayashi–Maskawa (CKM) phase is far too small to account for the observed asymmetry Shaposhnikov:1987tw ; Farrar:1993hn ; Gavela:1993ts ; Konstandin:2003dx ; Kapusta:2006pm . Consequently, achieving a strongly first-order transition and sufficient $CP$ violation for EWBG requires BSM physics. This mechanism has motivated extensive study of the EWPT and its realizations in BSM theories Cohen:1993nk ; Rubakov:1996vz ; Trodden:1998ym ; Riotto:1998bt ; Cline:2006ts ; Morrissey:2012db ; White:2016nbo ; Cline:2017jvp ; Wagner:2023vqw ; Liu:2023sey ; Goncalves:2023svb ; vandeVis:2025efm . Despite its appeal, no conclusive evidence for new electroweak-scale excitations has yet emerged. Furthermore, any additional $CP$-violating interactions introduced at the electroweak scale are strongly constrained by precision measurements of electric dipole moments (EDM) of the electron ACME:2018yjb , neutron Abel:2020pzs , and heavy atoms such as mercury Griffith:2009zz . A viable way to reconcile these limits with successful EWBG is through spontaneous $CP$ violation that occurs only during the EWPT. In this picture, the temporary breaking of $CP$ in the early universe supplies the baryogenesis source while the symmetry is restored at zero temperature, naturally evading the present EDM constraints.
@@ 2.1 The Inelastic Higgs-Portal Dark Matter Model
We add quartic self-interactions for the singlet fields $\phi_{1}$ and $\phi_{2}$, parametrized by $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{12}$. In the special case $\lambda_{1}=\lambda_{2}=\lambda_{12}$, the scalar sector retains an approximate $U(1)$ symmetry. More generally, allowing these couplings to differ explicitly breaks this symmetry. In the limits of $f_{1}\sim 0$ and $\sin 2\theta=0$, the coupling relations given in Eq. (13) to Eq. (15) and the mass relation are reduced to,
@@ 2.2.1 Quantum Corrections at Zero Temperature
One-loop corrections shift both the location of the electroweak vacuum and the scalar masses and mixings. To preserve the physical mass spectrum and vacuum expectation values at one loop, counterterms $V_{\rm CT}$ are introduced, parametrized by
@@ 2.2.2 Thermal Corrections
where the thermal functions $J_{B}$ and $J_{F}$ correspond to bosons and fermions, respectively:
@@ 3 Dark Matter Phenomenology
To assess the relative impact of the channels, we show $\Omega_{\phi_{1}}h^{2}$ versus $m_{\phi_{1}}$ for benchmark choices of $(\Delta m,\,g,\,f_{2})$ in Fig. 2. The coupling $f_{1}$ is fixed to a representative value $10^{-5}$, chosen sufficiently small to keep elastic Higgs-mediated scattering safely below current direct-detection bounds Hooper:2025fda ; LZCollaboration:2024lux . The horizontal dashed line marks the Planck value $\Omega_{\rm DM}h^{2}\simeq 0.12$ Planck:2018vyg ; vertical dotted lines at $m_{\phi_{1}}\simeq m_{h}/2$ and $m_{\phi_{1}}\simeq m_{h}$ indicate the Higgs resonance and the $hh$ threshold. Red/blue (green/black) curves correspond to $\Delta m=10$ GeV (0.1 GeV).
@@ 4 FOEWPT-favored parameter space
Figure 4: [Left]: Scan in the $m_{\phi_{2}}$$f_{2}$ plane showing points that realize an FOEWPT. We vary $m_{\phi_{2}}\in[100,250]~\mathrm{GeV}$ and $\Delta m\in[10,20]~\mathrm{GeV}$, while holding $\lambda_{1,2,12}\sim 0.5$ and $g\sim 0.05$ fixed. [Right]: Peak amplitude $\left(\Omega_{\mathrm{GW}}h^{2}\right)_{\mathrm{peak}}$ versus peak frequency $f_{\mathrm{peak}}$ of the predicted stochastic GW signal from FOEWPT-allowed points. Sensitivity curves for prospective GW observatories LISA, Taiji, ALIA, BBO, and UDECIGO are overlaid for comparison. In both panels, the color scale encodes the transition strength $\xi_{n}\equiv v_{n}/T_{n}$.
@@ 5 Spontaneous $CP$ violation and the generated baryon asymmetry
Benchmark point.- To illustrate our scenario, we specify a benchmark (Tab. 1) in which the DM mass is $\sim\!133~\mathrm{GeV}$ and the mass splitting is large enough that the relic density is set primarily by $\phi_{1}\phi_{1}\!\to hh$ via $\phi_{2}$ exchange annihilation channel. This choice yields a present-day di-Higgs indirect-detection cross section of $\langle\sigma v\rangle_{\rm ID}^{hh}\simeq 1.9\times 10^{-26}~\mathrm{cm^{3}\,s^{-1}}$, consistent with interpretations of the bright, statistically significant GeV excess from the Galactic Center for these masses Goodenough:2009gk ; Hooper:2010mq ; Hooper:2011ti ; Abazajian:2012pn ; Hooper:2013rwa ; Gordon:2013vta ; Daylan:2014rsa ; Calore:2014xka ; Zhou:2014lva ; Fermi-LAT:2015sau ; Fermi-LAT:2017opo ; Cholis:2021rpp ; DiMauro:2021raz ; Hooper:2019xss ; Hooper:2025fda ; Hu:2025thq . The elastic $\phi_{1}$$h$ coupling is chosen such that the spin–independent scattering rate ($\sigma_{\rm{DD}}^{\rm SI}$) lies below the “neutrino fog” Billard:2013qya .
This point also realizes the two-step thermal history central to our EWBG study. A second-order transition at $T_{n}=335~\mathrm{GeV}$ first breaks the $Z_{2}$ symmetry along the singlet directions so that $\langle\phi_{1,2}\rangle\neq 0$. At a lower temperature, $T_{n}=95~\mathrm{GeV}$, a first-order transition occurs in which electroweak symmetry is broken while the $Z_{2}$ is restored by driving the singlet $\it vevs$} back to zero. The field values at the false and true minima of the second transition are listed in Tab. 1. The corresponding bubble wall profiles $h(z)$, $\phi_{1}(z)$ and $\phi_{2}(z)$ are shown in the left plot of Fig. 5. As discussed earlier, it can be seen that all three fields remain nonzero in certain region (within the bubble wall region), where the $CP$-violating source built. In this benchmark scenario $L_{w}T_{n}=3.5$ and we set $v_{w}=0.1$. The generated BAU tracks the effective left-handed baryon chemical potential $\mu_{B_{L}}(z)$ defined in Eq. (76), whose profile for this point is shown in the right plot of Fig. 5. To match the observed value of BAU, we set $c=1$ and vary the cutoff scale and fix it at $\Lambda\simeq 800~\mathrm{GeV}$ for this benchmark point. A deep scan in the surrounding parameter space can be performed, which is expected to preserve the qualitative features of the benchmark while achieving the observed BAU with somewhat larger values of $\Lambda$.
@@ Acknowledgments
The author gratefully acknowledges Dan Hooper, Gordan Krnjaic, Tong Ou and Carlos Wagner for many insightful discussions. This work is supported by the U.S. Department of Energy under contract No. DEAC02-06CH11357 at Argonne National Laboratory. The author further acknowledges the hospitality of the University of Chicago, Fermilab, the Aspen Center for Physics, and the Perimeter Institute for Theoretical Physics, where a significant part of this research was carried out. He is especially indebted to Carlos Wagner for hosting him at Perimeter Institute.